Filter  Source  Sky (e-/second) (e-/second/pixel) U  32  1.4 B  164  3.1 V  181  10.5 R  151  11.5 I  113  14.6 Rough exposure times to reach

mag = 22 (delta mag= 0.05m) in the five filters are: U 1250s B 120s V 320s R 500s I 1085s